# MC-GPT: Empowering Vision-and-Language Navigation with Memory Map and Reasoning Chains

### v1

@@ Abstract.
vision and language navigation, neural networks, prompt-based learning
copyright: noneccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Vision for robotics
Figure 1. Accurate memory and suitable strategies are crucial for navigation tasks. A topological map can precisely record information about viewpoints, while incorporating human navigation reasoning examples can enhance the diversity of strategies employed by the LLM.
1. Introduction
In recent years, the convergence of computer vision and natural language processing has given rise to an emerging task known as Vision-and-Language Navigation (VLN) (anderson2018vision, ; qi2020reverie, ; ku2020room, ). This cross-modal task necessitates the integration of visual perception and natural language understanding, enabling autonomous agents to understand their surrounding environments and navigate to destinations. The VLN task requires agents to follow natural language instructions and align content from textual and visual modalities.
@@ 1. Introduction
Recently, Large Language Models (LLMs) (brown2020language, ; driess2023palm, ; touvron2023llama, ) have emerged as a promising approach to addressing the issues with VLN tasks. The generalization ability of LLMs can help mitigate overfitting issues and reduce training costs, while the common sense knowledge of LLMs can provide enhanced interpretability for navigation reasoning. Some efforts have integrated LLMs as sub-modules within traditional navigation frameworks, allowing LLMs to handle subtasks such as instruction decomposition and landmark extraction (shah2023lm, ; zhou2023esc, ; qiao2023march, ; cai2023bridging, ). We refer to this as the LLM-sub approach. With this approach, the overfitting issue persists as it still requires training. Conversely, a few studies have positioned LLMs as core controllers in navigation decision-making (zhou2023navgpt, ; chen2024mapgpt, ; dorbala2023can, ; schumann2023velma, ) without any additional training, an approach we call LLM-core. In this paper, we adopt the LLM-core approach in order to fully exploit the potential of the LLM and maximize the generalization ability of our method.
@@
Despite the advancements, existing LLM-core methods tend to oversimplify problem modeling. For instance, most methods utilize a combination of Vision-and-Language Model (VLM) (li-etal-2023-lavis, ; liu2024llava, ; achiam2023gpt4, ) and LLM to summarize the navigation history as a context, or memory, for LLM interaction. This approach not only introduces noise but also requires multiple calls of models, which increases time consumption and API costs. It is suggested that the key information that affects VLN performance includes objects, room type, and orientation (qi2021know, ; gao2021room, ). Therefore, we propose constructing a memory structure that accurately captures the navigation context, and highlights these components. As depicted in Fig. 1, we introduce a topological map as the memory structure. The map records the information about viewpoints and objects, as well as their spatial relationships. During navigation, the agent adds observations as new nodes to the topological map, and updates the layout of objects accordingly. The connectivity between nodes preserves the orientation information, while the object layouts retains object information, and allows for inferring room-type information.
@@
Typically, the current LLM-core methods need to provide basic rules for navigation, a technique known as Standard Prompting (brown2020language, ). However, prescribing the reasoning process can limit the agent’s navigation strategies, and has difficulty incorporating domain-specific knowledge. In reality, humans may employ different navigation strategies for different scenarios. For example, executing the instructions “go to the kitchen to get some milk” and “go to the utility room to get a stool” requires distinct navigation strategies. Here, “kitchen” is an easily recognizable room type, which provides clear directional guidance. In contrast, the utility room may be less common, with a somewhat ambiguous definition, as it may coexist with other areas of a house, such as the laundry room. In essence, executing the first task may involve a simple room-type based priority strategy, while the second task requires a strategy that balances room-type priority with object priority, alongside commonsense reasoning.
@@
The reasoning process on such a case-by-case basis can be addressed by having the LLM look at demonstration examples (brown2020language, ; min2022rethinking, ). The benefit is more pronounced when the offered examples contain a step-by-step reasoning process, a method called the Chain-of-Thought (CoT) prompting (wei2022chain, ; zhang2022automatic, ). Inspired by these techniques, we introduce a Navigation Chain of Thought Module, as depicted in Fig. 1. We first establish an example set covering various scenarios, and employ an LLM to mine the reasoning chains of these examples; during testing, the LLM can retrieve a demonstration example from this set. Notably, the examples are collected from navigation datasets such as REVERIE (qi2020reverie, ). These datasets are annotated by humans and thus reflect human navigation decision-making. Our approach of navigation does not need extra manual annotation.
@@
2. Related work
2.1. Vision-and-Language Navigation
@@
The Vision-and-Language Navigation (VLN) task (anderson2018vision, ; qi2020reverie, ; ku2020room, ) requires robots to understand natural language navigation instructions and move to target locations by executing a series of navigation actions. Current VLN methods primarily focus on improving navigation success rates and efficiency by optimizing model structure (anderson2018vision, ; hong2021vln, ; chen2022think, ; qiao2022hop, ; liang2022visual, ; lin2022adapt, ; an2022bevbert, ), representation learning(wang2020soft, ; wang2019reinforced, ), and data augmentation(fu2020counterfactual, ; zhan2023object, ).
@@
To address the issue of data scarcity, Fu et al. (fu2020counterfactual, ) utilize counterfactual reasoning for data augmentation. They make the path sampler and navigator against each other during training based on the principles of adversarial learning. Wang et al. (wang2020soft, ) improve supervised signals by introducing additional terms to the loss function of the reinforcement learning based on knowledge distillation, making the supervision signals softer and more robust. Wang et al. (wang2019reinforced, ) find that training the navigator with the shortest path as the sole supervision signal may lead to biases in training when human annotations do not correspond strictly with the shortest path. Therefore, using the distance to the navigation endpoint as an additional supervision signal can enhance training robustness. This training strategy can also be generalized to synthetic data. With the rise of pre-trained models and the development of prompting engineering, a series of VLN methods (hong2021vln, ; chen2022think, ; qiao2022hop, ; liang2022visual, ; lin2022adapt, ; an2022bevbert, ) based on pre-trained models have been proposed. These methods improve the effectiveness of VLN algorithms in various ways, although they may result in increased costs in computation and training time.
2.2. Navigation with LLM
@@
LLM’s generalization and reasoning capabilities have sparked interest in navigation tasks (brown2020language, ; driess2023palm, ; touvron2023llama, ). Some researchers attempt to use the LLM as an auxiliary module in the traditional navigation frameworks, in which the LLM provides navigation guidance or helps with landmark extraction. Shah et al. (shah2023lm, ) utilize the GPT-3 to identify landmarks in instructions, aligning textual and visual modalities using a VLM (CLIP (radford2021learning, )), and then integrate them into a traditional navigation framework to complete navigation reasoning. Zhou et al. (zhou2023esc, ) introduce VLM as a scene understanding module, leveraging large language models to provide commonsense knowledge, using probabilistic soft logic to formulate reasoning rules, and finally combining them with frontier exploration strategies to predict navigation actions. Qiao et al. (qiao2023march, ) focus on high-level concise navigation instructions, employing large language models to generate fine-grained instructions. They predict target objects, destinations, and current objects and room types, and when a change in room type is perceived, they use the LLM to generate fine-grained next-step instruction. Cai et al. (cai2023bridging, ) utilize the GPT-4 (achiam2023gpt4, ) for self-localization and next-step decision-making, then use a pixel-based navigation algorithm to complete subsequent navigation actions.
@@
In other works, LLM is employed as the core controller of navigation, where the LLM directly makes navigation decisions and interacts with other modules within the navigation system. Zhou et al. (zhou2023navgpt, ) describe visual perception using VLM, detect objects using an object detector, and then employ the ChatGPT to summarize historical information and make specific reasoning. Chen et al. (chen2024mapgpt, ) introduce a global map into an LLM-based navigator, expanding the navigator’s action space from local to global. Dorbala et al. (dorbala2023can, ) use the LLM for navigation decision-making, directing navigation towards the target direction when no target objects are perceived, and navigating around the target object when it appears in sight. Schumann et al. (schumann2023velma, ) propose an embodied LLM agent, using language descriptions of trajectories and visual environment observations as context hints for the next action, and outputting navigation actions based on LLM’s next-word prediction. Different from previous works, our method aims to address the limitations of LLM-based navigators in memory construction and the diversity of navigation strategies. We also propose a novel pipeline for LLM-based navigation.
3. Approach
3.1. Problem setup
@@
According to the setting of the VLN task, an agent is randomly placed at a viewpoint and receives an instruction. To carry out the instruction, the agent needs to reason step-by-step by grounding the textual instruction and visual observation. At each time step, the agent selects and moves to one of the navigable viewpoints according to the instruction and the observation. Under the REVERIE task (object-oriented VLN task) setting, the agent also needs to point out the referred object from a set of object proposals.
3.2. Overview
@@
Fig. 2 presents our pipeline. Initially, our agent acquires visual observations of the surrounding environments and stores them in memory using the Memory Topological Map module. Further elaboration on the Memory Topological Map module can be found in Section 3.3. Subsequently, our agent utilizes the Navigation Chain of Thoughts module to query an example with a chain of thoughts by computing the similarities between the test instance and the example set. More details regarding the Navigation Chain of Thoughts module are provided in Section 3.4. Finally, the agent organizes the information, including Task Description, Instruction, Memory Topological Map, and Demonstration Example, using the Prompt Manager, and passes them to the LLM. Details of the Prompt Manager can be found in Section 3.5. The LLM then conducts reasoning and decides which viewpoint to select. Subsequently, the agent employs a Shortest-Path Planner to move to the selected viewpoint.
@@
Figure 2. Overview of our pipeline. The Memory Topological Map module manages a map for storing memory, while the Navigation Chain of Thoughts module provides a demonstration example with a reasoning chain. The Prompt manager organizes all essential information and feeds it to the LLM. The LLM then makes a viewpoint selection and employs a Shortest-path Planner to navigate to the selected viewpoint.
3.3. Memory Topological Map module
@@
The Memory Topological Map records the information of observed viewpoints (including visited and unvisited ones) in a topological graph. The construction process is illustrated in Fig. 3 and can be roughly divided into five steps as follows.
@@
Figure 3. The construction of the Memory Topological Map. Initially, with object detection and image captioning, we can convert visual perceptions into textual descriptions. Subsequently, we construct a map comprising viewpoints and objects. Notably, the red node represents the viewpoint where the agent is currently located, while the yellow nodes denote unvisited viewpoints, and the blue nodes mean visited viewpoints. The squares on the map represent objects, and squares with the same colors indicate that they are clustered together. Finally, we utilize a textual prompt template to interpret the map.
@@
Object Detection. For each viewpoint, we retain its object detection results. Specifically, for newly observed viewpoints, we capture images from the current viewpoint and send them to the Faster R-CNN with an attention mechanism (anderson2018bottom, ) for object detection. For existing viewpoints, we persist with the previously obtained object detection results.
@@
Image Captioning. To enhance comprehension of the visual content within viewpoints, we utilize a VLM named InstructBLIP (li-etal-2023-lavis, ) to generate captions for their images. To minimize computational costs, we selectively invoke InstructBLIP for neighboring viewpoints only, while retaining previously generated captions for other viewpoints.
@@
Mapping. We then create a topological map incorporating object distribution. Objects detected solely in one viewpoint are positioned at that respective viewpoint. However, for objects observed across multiple viewpoints, we position them at the geometric center of the corresponding viewpoints. Besides, considering the co-occurrence of objects provides insight into the room type, we employ the K-Means clustering algorithm (macqueen1967some, ; hartigan1979algorithm, ) to cluster nearby objects together. The resulting object-clustered outcomes are utilized for room-type prediction, which is a sub-task of our navigation reasoning. The reason for the sub-task is that the room-type priority strategy is an important navigation strategy for those recognizable room types.
@@
Interpreting. In order to facilitate the understanding of the map by the LLM, textual interpretation is necessary. Drawing inspiration from (vemprala2023chatgpt, ), we adopt a structured language template: ”Memory topological map: —cluster ID: objects—, Navigable viewpoints: ¡viewpoint ID, image captioning, relationship with Object clusters¿”. As shown in the top-right of Fig 5, the Memory topological map field provides the object information and the clusters, while the Navigable viewpoints field provides the connection relationship between the viewpoints and clusters. Following interpretation, the map is converted into a textual format, which can be directly provided to the LLM as part of the prompt.
3.4. Navigation Chain of Thoughts module

### v2

@@ Abstract
Index Terms:
vision and language processing, navigation, large language model, multi-modal large language model.
Figure 1: Accurate memory and suitable strategies are crucial for navigation tasks. A topological map can precisely record information about viewpoints, while incorporating human navigation reasoning examples can enhance the diversity of strategies employed by the LLM.
I Introduction
In recent years, the convergence of computer vision and natural language processing has given rise to an emerging task known as Vision-and-Language Navigation (VLN) [1, 2, 3]. This cross-modal task necessitates the integration of visual perception and natural language understanding, enabling autonomous agents to understand their surrounding environments and navigate to destinations. The VLN task requires an agent to follow a natural language instruction and navigate to the destination.
@@ I Introduction
Recently, Large Language Models (LLMs) [16, 17, 18] have emerged as a promising approach to addressing the issues with VLN tasks. The generalization ability of LLMs can help mitigate overfitting issues and reduce training costs, while the common sense knowledge of LLMs can provide enhanced interpretability for navigation reasoning. Some efforts have integrated LLMs as sub-modules within traditional navigation frameworks, allowing LLMs to handle subtasks such as instruction decomposition and landmark extraction [19, 20, 21, 22]. We refer to this as the LLM-sub approach. With this approach, the overfitting issue persists as it still requires training. Conversely, a few studies have positioned LLMs as core controllers in navigation decision-making [23, 25, 26, 27] without any additional training, an approach we call LLM-core. In this paper, we adopt the LLM-core approach in order to fully exploit the potential of the LLM and maximize the generalization ability of our method.
@@
Despite the advancements, existing LLM-core methods tend to oversimplify problem modeling. For instance, most methods utilize a combination of Vision-and-Language Model (VLM) [29, 31, 30, 47] and LLM to summarize the navigation history as a context, or memory, for LLM interaction. This approach not only introduces noise but also requires multiple calls of models, which increases time consumption and API costs. It is suggested that the key information that affects VLN performance includes objects, room type, and orientation [32, 33]. Therefore, we propose constructing a memory structure that accurately captures the navigation context, and highlights these components. As depicted in Fig. 1, we introduce a topological map as the memory structure. The map records the information about viewpoints and objects, as well as their spatial relationships. During navigation, the agent adds observations as new nodes to the topological map, and updates the layout of objects accordingly. The connectivity between nodes preserves the orientation information, while the object layouts retains object information, and allows for inferring room-type information.
@@
Typically, the current LLM-core methods need to provide basic rules for navigation, a technique known as Standard Prompting [16]. However, prescribing the reasoning process can limit the agent’s navigation strategies, and has difficulty incorporating domain-specific knowledge. In reality, humans may employ different navigation strategies for different scenarios. For example, executing the instructions “go to the kitchen to get some milk” and “go to the utility room to get a stool” requires distinct navigation strategies. Here, “kitchen” is an easily recognizable room type, which provides clear directional guidance. In contrast, the utility room may be less common, with a somewhat ambiguous definition, as it may coexist with other areas of a house, such as the laundry room. In essence, executing the first task may involve a simple room-type based priority strategy, while the second task requires a strategy that balances room-type priority with object priority, alongside commonsense reasoning.
@@
The reasoning process on such a case-by-case basis can be addressed by having the LLM look at demonstration examples [16, 35]. The benefit is more pronounced when the offered examples contain a step-by-step reasoning process, a method called the Chain-of-Thought (CoT) prompting [36, 37]. Inspired by these techniques, we introduce a Navigation Chain of Thought Module, as depicted in Fig. 1. We first establish an example set covering various scenarios, and employ an LLM to mine the reasoning chains of these examples; during testing, the LLM can retrieve a demonstration example from this set. Notably, the examples are collected from navigation datasets such as REVERIE [2]. These datasets are annotated by humans and thus reflect human navigation decision-making. Our approach of navigation does not need extra manual annotation.
@@
II Related work
II-A Vision-and-Language Navigation
@@
The Vision-and-Language Navigation (VLN) task [1, 2, 3] requires robots to understand natural language navigation instructions and move to target locations by executing a series of navigation actions. Current VLN methods primarily focus on improving navigation success rates and efficiency by optimizing model structure [1, 14, 10, 15, 11, 12, 13], representation learning[5, 6], and data augmentation[4, 28].
@@
To address the issue of data scarcity, Fu et al. [4] utilize counterfactual reasoning for data augmentation. They make the path sampler and navigator against each other during training based on the principles of adversarial learning. Wang et al. [5] improve supervised signals by introducing additional terms to the loss function of the reinforcement learning based on knowledge distillation, making the supervision signals softer and more robust. Wang et al. [6] find that training the navigator with the shortest path as the sole supervision signal may lead to biases in training when human annotations do not correspond strictly with the shortest path. Therefore, using the distance to the navigation endpoint as an additional supervision signal can enhance training robustness. This training strategy can also be generalized to synthetic data. With the rise of pre-trained models and the development of prompting engineering, a series of VLN methods [14, 10, 15, 11, 12, 13] based on pre-trained models have been proposed. These methods improve the effectiveness of VLN algorithms in various ways, although they may result in increased costs in computation and training time.
II-B Navigation with LLM
@@
LLM’s generalization and reasoning capabilities have sparked interest in navigation tasks [16, 17, 18, 24]. Some researchers attempt to use the LLM as an auxiliary module in the traditional navigation frameworks, in which the LLM provides navigation guidance or helps with landmark extraction. Shah et al. [19] utilize the GPT-3 to identify landmarks in instructions, aligning textual and visual modalities using a VLM (CLIP [34]), and then integrate them into a traditional navigation framework to complete navigation reasoning. Zhou et al. [20] introduce VLM as a scene understanding module, leveraging large language models to provide commonsense knowledge, using probabilistic soft logic to formulate reasoning rules, and finally combining them with frontier exploration strategies to predict navigation actions. Qiao et al. [21] focus on high-level concise navigation instructions, employing large language models to generate fine-grained instructions. They predict target objects, destinations, and current objects and room types, and when a change in room type is perceived, they use the LLM to generate fine-grained next-step instruction. Cai et al. [22] utilize the GPT-4 [30] for self-localization and next-step decision-making, then use a pixel-based navigation algorithm to complete subsequent navigation actions.
@@
In other works, LLM is employed as the core controller of navigation, where the LLM directly makes navigation decisions and interacts with other modules within the navigation system. Zhou et al. [23] describe visual perception using VLM, detect objects using an object detector, and then employ the ChatGPT to summarize historical information and make specific reasoning. Chen et al. [25] introduce a global map into an LLM-based navigator, expanding the navigator’s action space from local to global. Dorbala et al. [26] use the LLM for navigation decision-making, directing navigation towards the target direction when no target objects are perceived, and navigating around the target object when it appears in sight. Schumann et al. [27] propose an embodied LLM agent, using language descriptions of trajectories and visual environment observations as context hints for the next action, and outputting navigation actions based on LLM’s next-word prediction. Different from previous works, our method aims to address the limitations of LLM-based navigators in memory construction and the diversity of navigation strategies. We also propose a novel pipeline for LLM-based navigation.
III Approach
III-A Problem setup
@@
According to the setting of the VLN task, an agent is randomly placed at a viewpoint and receives an instruction. To carry out the instruction, the agent needs to reason step-by-step by grounding the textual instruction and visual observation. At each time step, the agent selects and moves to one of the navigable viewpoints according to the instruction and the observation. Under the REVERIE task (object-oriented VLN task) setting, the agent also needs to carry out the referring expression task [42] - point out the referred object from a set of object proposals.
Formally, let I be the navigation instruction, and we can obtain the visual observation by calling the function of MP3D simulator [1]:
${V_{t}}{=\left\{{{\emph{{v}}_{j}}}\right\}_{j=1}^{J}=\rm{get\_observation()}}$ (1)
where J is the number of discrete views. Among J discrete views, only L views could be reached, we call them local navigable viewpoints and denote them as $\hat{\emph{{{V}}}}=\left\{{\emph{{v}}_{l}}\right\}_{l=1}^{L}$, where $L\leq J$. By combining the previous local navigable viewpoints and a “STOP” state, we can formulate the global navigational viewpoints a.k.a action space:
${\emph{{U}}_{t}}=\left\{{{\emph{{u}}_{n}}}\right\}_{n=1}^{N}=\sum_{h=0}^{t}\hat{\emph{{{V}}}}_{h}\cup\left\{STOP\right\}$ (2)
$=\rm{get\_navigational\_viewpoints}(\emph{V}_{t})$
where $N$ is the number of global navigable viewpoints. At each time step, the agent selects and moves to one of the navigable viewpoints ${\emph{{U}}_{t}}$. Under the object-oriented navigation task setting such as the REVERIE task, the agent also needs to point out the referred object from a set of object proposals:
$\emph{{O}}_{t}=\left\{{{\emph{{o}}_{k}}}\right\}_{k=1}^{K}=\rm{get\_object\_proposals}(\emph{V}_{t})$ (3)
where K is the number of object proposals.
III-B Overview
@@
Fig. 2 and Algorithm 1 present our pipeline. Initially, our agent acquires visual observations of the surrounding environments and stores them in memory using the Memory Topological Map module. Further elaboration on the Memory Topological Map module can be found in Section III-C. Subsequently, our agent utilizes the Navigation Chain of Thoughts module to query an example with a chain of thoughts by computing the similarities between the test instance and the example set. More details regarding the Navigation Chain of Thoughts module are provided in Section III-D. Finally, the agent organizes the information, including Task Description, Instruction, Memory Topological Map, and Demonstration Example, using the Prompt Manager, and passes them to the LLM. Details of the Prompt Manager can be found in Section III-E. The LLM then conducts reasoning and decides which viewpoint to select. Subsequently, the agent employs a Shortest-Path Planner to move to the selected viewpoint.
@@
Figure 2: Overview of our pipeline. The Memory Topological Map module manages a map for storing memory, while the Navigation Chain of Thoughts module provides a demonstration example with a reasoning chain. The Prompt manager organizes all essential information and feeds it to the LLM. The LLM then makes a viewpoint selection and employs a Shortest-path Planner to navigate to the selected viewpoint.
Algorithm 1 Task Execution Algorithm
1:  Input: Task description $T$, instruction $I$, example set $D=\{d_{1},d_{2},\ldots,d_{Q}\}$, maximum number of navigation steps $max\_episode\_length$
2:  Output: Selected viewpoint $u_{n}$, selected object $o_{k}$
3:  Operator functions:
4:  $\rm{F_{M}}$: Memory Topological Map module
5:  $\rm{F_{C}}$: Navigation Chain of Thoughts module
6:  $\rm{F_{P}}$: Prompt manager
7:  $\rm{LLM}$: Large language model
8:  $\rm{F_{S}}$: Shortest-path planner
9:  $step\leftarrow 0$
10:  for $step<max\_episode\_length$ do
11:     // Obtain observation, navigational viewpoints and object proposals from simulator
12:     $V_{t}\leftarrow get\_observation()$
13:     $U_{t}\leftarrow get\_navigational\_viewpoints(V_{t})$
14:     $O_{t}\leftarrow get\_object\_proposals(V_{t})$
15:     // Construct memory topological map
16:     $M_{t}\leftarrow\rm{F_{M}}(\emph{U}_{t},\emph{O}_{t},\emph{M}_{t-1})$
17:     // Query a chain of thought example
[…]
19:     // Build prompt
20:     $P_{t}\leftarrow\rm{F_{P}}(\emph{M}_{t},\emph{T},\emph{I},\emph{d}_{q})$
21:     // Select a navigational viewpoint and object
22:     $(u_{n}\in U_{t},o_{k}\in O_{t})\leftarrow\rm{LLM}(\emph{P}_{t})$
23:     // Execute action
24:     $\rm{F_{S}}(\emph{u}_{n},\emph{M}_{t})$
25:     if $u_{n}==\rm{STOP}$ then
26:        break
27:     end if
28:     $step\leftarrow step+1$
29:  end for
30:  return  $u_{n},o_{k}$
III-C Memory Topological Map module
@@
The Memory Topological Map records the information of observed viewpoints (including visited and unvisited ones) in a topological graph. Formally, the construction of a memory topological map can be represented as:
$M_{t}=\rm{F_{M}}(\emph{U}_{t},\emph{O}_{t},\emph{M}_{t-1})$ (4)
The operations of $\rm{F_{M}}$ are illustrated in Fig. 3 and can be roughly divided into five steps as follows.
@@
Figure 3: The construction of the Memory Topological Map. Initially, with object detection and image captioning, we can convert visual perceptions into textual descriptions. Subsequently, we construct a map comprising viewpoints and objects. Notably, the red node represents the viewpoint where the agent is currently located, while the yellow nodes denote unvisited viewpoints, and the blue nodes mean visited viewpoints. The squares on the map represent objects, and squares with the same colors indicate that they are clustered together. Finally, we utilize a textual prompt template to interpret the map.
@@
Object Detection. For each viewpoint, we retain its object detection results. Specifically, for newly observed viewpoints, we capture images from the current viewpoint and send them to the Faster R-CNN with an attention mechanism [38] for object detection. For existing viewpoints, we persist with the previously obtained object detection results.
@@
Image Captioning. To enhance comprehension of the visual content within viewpoints, we utilize a VLM named InstructBLIP [29] to generate captions for their images. To minimize computational costs, we selectively invoke InstructBLIP for neighboring viewpoints only, while retaining previously generated captions for other viewpoints. Some qualitative examples can be found in Sec. V-D.
@@
Mapping. We then create a topological map incorporating object distribution. Objects detected solely in one viewpoint are positioned at that respective viewpoint. However, for objects observed across multiple viewpoints, we position them at the geometric center of the corresponding viewpoints. Besides, considering the co-occurrence of objects provides insight into the room type, we employ the K-Means clustering algorithm [39, 40] to cluster nearby objects together. The resulting object-clustered outcomes are utilized for room-type prediction, which is a sub-task of our navigation reasoning. The reason for the sub-task is that the room-type priority strategy is an important navigation strategy for those recognizable room types.
@@
Interpreting. In order to facilitate the understanding of the map by the LLM, textual interpretation is necessary. Drawing inspiration from [41], we adopt a structured language template: “Memory topological map: ||||cluster ID: {objects}||||, Navigable viewpoints: <<<viewpoint ID, image captioning, relationship with Object clusters>>>”. As shown in the top-right of Fig 5, the Memory topological map field provides the object information and the clusters, while the Navigable viewpoints field provides the connection relationship between the viewpoints and clusters. Following interpretation, the map is converted into a textual format, which can be directly provided to the LLM as part of the prompt.
Algorithm 2 CoT Mining Algorithm
1:  Input: Task description $T^{\prime}$, instruction $I$, sampled training set $S$
2:  Output: Example set $D$
3:  Operator functions:
4:  $\rm{F_{M}}$: Memory Topological Map module
5:  $\rm{F_{P}}$: Prompt manager
6:  $\rm{LLM}$: Large language model
7:  $\rm{F_{S}}$: Shortest-path planner
8:  // Initialize an example set
9:  $D\leftarrow[]$
10:  for $(i,ground\_truth\_path)\in\ S$ do
11:     // Initialize a CoT example
12:     $d\leftarrow[]$
13:     for $(u_{n},o_{k})\in\ ground\_truth\_path$ do
14:        // Obtain observation, navigational viewpoints and object proposals from simulator
15:        $V_{t}\leftarrow get\_observation()$
16:        $U_{t}\leftarrow get\_navigational\_viewpoints(V_{t})$
17:        $O_{t}\leftarrow get\_object\_proposals(V_{t})$
18:        // Construct memory topological map
19:        $M_{t}\leftarrow\rm{F_{M}}(\emph{U}_{t},\emph{O}_{t},\emph{M}_{t-1})$
20:        // Build prompt
21:        $P_{t}\leftarrow\rm{F_{P}}(\emph{M}_{t},\emph{T'},\emph{I},\emph{u}_{n},\emph{o}_{k})$
22:        // Explain the reason behind viewpoint and object selection
23:        $explanation\leftarrow\rm{LLM}(\emph{P}_{t})$
24:        // Update the CoT example with a “instruction-path-explanation” pair
25:        Add $(I,u_{n},o_{k},explanation)$ to $d$
26:        // Execute action
27:        $\rm{F_{S}}(\emph{u}_{n},\emph{M}_{t})$
28:        if $u_{n}==\rm{STOP}$ then
29:           break
30:        end if
31:     end for
32:     // Update the example set
33:     Add $d$ to $D$
34:  end for
35:  return  $D$
III-D Navigation Chain of Thoughts module